OpenStudy (anonymous):

(2^5)^3x2^6 A. 2^48 B. 2^21 C. 2^9 D. 2^14

OpenStudy (anonymous):

B

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

(2^5)^3 = 2^ 15 and 2^15+2^6 = 2^21